# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ 2.2 Statements of the Main Results
In the case of subquintic sources, the analysis is simpler [S-S solutions are global], though still technical. Some of the recent relevant references include [24, 28] and references within, where nonlinear dissipation has been treated. For the quintic waves with nonlinear dissipation some of the elements of the S-S method described above has been very recently applied to subquintic and quintic waves with nonlinear enhanced dissipation [30]. In particular, backward trajectories method patterned after [29] is the driving argument in establishing asymptotic compactness of the trajectories and their smoothness. This requires strong assumptions imposed on the damping [nondegenearcy of the slope at the origin] and also the source [coercivity of quintic estimate].
@@
One oof the goals of this paper is to show that such restrictions, in general, are not needed. The model with enhanced dissipation takes its own life with no reference to either S-S solutions or backward trajectories. The main protagonist is the enhanced dissipation which could also be degenerate $g^{\prime}(0)=0$.] This alone provides asymptotic compactness effect for a large range of sources including the quintic. No coercivity assumption is needed. The role of Strichartz estimates is taken, in some sense, by the enhanced dissipation generated by the nonlinearity of the damping. However, in the present case one is in a position to show that the enhanced dissipation is retained in the ultimate dissipativity estimate [unlike Strichartz estimate]. In view of this, the present work is a continuation of the model considered [5] in the direction of dynamical systems and its long time behaviour. The main player in this game is energy equality established for weak solutions - an earlier version of the equality can be found in [5]. Adaptation of J.Ball’s method to nonlinear dissipation leads to the existence of global attractor. A general quasi-stability property for the dynamics is established under additional assumption of nonvanishing derivative of the damper. The latter implies finite-dimensionality, enhanced regularity of the attractor and also an existence of exponential attractor. In order to recover maximal regularity of the attractor, the splitting method used by Chueshov and Lasiecka in [12] in conjunction with gradient structure of the attractor and properties of the stationary set. In conclusion, the results provided eliminate critical restrictions in [30] and moreover lead to different methodology applicable to many other hyperbolic like dynamics with critical sources. The method is much simpler and it could also be applied to nonlocal damping under certain assumptions [30, 31].
@@
One should also mention works where strong damping is added to the model. But this is a form of regularization where originally hyperbolic dynamics become parabolic [with linearisation generating an analytic semigroup]. The latter is a very different problem. References [10, 11, 17, 19].
The reminder of this paper is devoted to the proofs of the three theorems. Appendix 7 at the end of the paper, collects various definitions and well known facts used in the manuscript.
@@
3 Proof of Theorem 2.2: Weak Solutions
@@
3.1 Proof of Theorem 2.2-(I): Construction of Weak Solutions
The proof of existence of weak solutions proceeds along general strategy used in [5], where the enhanced dissipation plays critical role in the arguments. However, the construction in [5] is based on suitable truncations of nonlinear terms and heavy use of monotone operator theory [3, 22]. In the present case, we resort to Galerkin solutions, which also give a nice computational framework. The advantage of Galerkin solutions is that “formal” computations are well justified and lead to conserved quantities. However, critical limit passages on nonlinear terms rely on some arguments seen already in [5]. On the other hand, [5] deals also with the boundary sources of critical range.These will be more difficult to incorporate within the Galerkin scheme.
@@
Let $0<\lambda_{1}\leq\lambda_{2}\leq\cdots$ be the eigenvalues of the Laplacian operator $\Delta$ with Dirichlet boundary condition and $\omega_{1},\omega_{2},\cdots$ be the corresponding eigenfunctions such that they form an orthonormal basis in $H^{0}$. We assume the boundary $\Gamma=\partial\Omega$ regular enough such that $\omega_{j}\in H^{2}.$ Let $V_{k}=\mbox{Span}\{\omega_{1},\cdots,\omega_{k}\}$ the subspace of $H^{2}$ generated by the first $k$ elements of $\{\omega_{j}\}_{j\in\mathbb{N}}$. For $k\in\mathbb{N}$ we can construct a function $u^{k}$ given by
$u^{k}(t)=\sum_{j=1}^{k}y_{jk}(t)\,\omega_{j}\in V_{k},\quad t\in[0,t_{k}),$
where $(y_{jk})$ is a local solution on $0,t_{k})\in[0,T)$ of the following system of ODEs:
$\left\{\begin{array}[]{l}\displaystyle{(u^{k}_{tt}(t),\omega_{j})+(\nabla u^{k}(t),\nabla\omega_{j})+(g(u^{k}_{t}(t)),\omega_{j})+(f(u^{k}(t)),\omega_{j})=(h,\omega_{j})}\\ \displaystyle{U^{k}(0)=(u^{k}(0),u^{k}_{t}(0))=(u_{0k},u_{1k})=U_{0k},\quad j=1,\cdots,k.}\end{array}\right$. (3.3)
@@
$U^{k}(0)\longrightarrow U_{0}\quad\mbox{strongly in}\quad\mathcal{H}.$ (3.4)
Taking $\omega_{j}=u^{k}_{t}(t)$ in the approximate equation (3.3) and integrating over $[0,t]$ yields
$E(U^{k}(t))+\int_{0}^{t}g(u^{k}_{t})u^{k}_{t}dxds=E(U^{k}(0)),$ (3.5)
where $E$ is given in (2.9). Next, from (2.5) and embedding $H^{1}\hookrightarrow H^{0}$, we get
$\int_{\Omega}F(u^{k})dx\geq-\frac{\nu}{2}\|u^{k}(t)\|^{2}-C_{\nu}|\Omega|\geq-\frac{\nu}{2\lambda_{1}}\|\nabla u^{k}(t)\|^{2}-C_{\nu}|\Omega|.$ (3.6)
And, from Hölder inequality, embedding $H^{1}\hookrightarrow H^{0}$, and Young’s inequality, we have
$\int_{\Omega}hu^{k}dx\leq\|h\|\|u^{k}(t)\|\leq\frac{1}{\lambda_{1}^{1/2}}\|h\|\|\nabla u^{k}(t)\|\leq\frac{1}{\delta\lambda_{1}}\|h\|^{2}+\frac{\delta}{4}\|\nabla u^{k}(t)\|^{2},$ (3.7)
for any $\delta>0$. Then, choosing in particular $\delta=\omega$ (as defined in (2.6)), combining estimates (3.6) and (3.7), we can estimate the energy $E(U^{k})$ lower as follows
[…]
Using (2.1), (3.8), and convergence (3.4), it follows from (3.5) that
[…]
in $0,t_{k})$, $\forall k\in\mathbb{N}$. Now let’s estimate the term $E(U_{0k})$. Indeed, from (2.4)-(2.5), we have
$\int_{\Omega}F(u_{0k})dx$ $\leq$ $\int_{\Omega}f(u_{0k})u_{0k}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$=$ $\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u_{0k})d\theta u_{0k}^{2}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$\leq$ $C_{f}\int_{\Omega}\int_{0}^{1}(1+|\theta u_{0k}|^{4})d\theta u_{0k}^{2}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$\leq$ $\left(C_{f}+\frac{\nu}{2}\right)\|u_{0k}\|^{2}+C_{f}\|u_{0k}\|^{6}_{6}.$
Hence, using embeddings $H^{1}\hookrightarrow L^{6}(\Omega)\hookrightarrow H^{0}$ and convergence (3.4), we have
$E(U_{0\kappa})\leq\frac{1}{2}\|u_{1k}\|^{2}+\frac{1}{2}\|\nabla u_{0k}\|^{2}+\left(C_{f}+\frac{\nu}{2}\right)\|u_{0k}\|^{2}+C_{f}\|u_{0k}\|^{6}_{6}+\|h\|\|u_{0k}\|\leq C,$
@@
for $t\in[0,t_{k})$. Therefore, estimate (3.10) allows us to extend the local solution of the approximate problem to whole interval $[0,T]$, for any given $T>0$. From (3.10), we have
$U^{k}=(u^{k},u^{k}_{t})\quad$ is bounded in $\quad L^{\infty}(0,T;\mathcal{H}),$ (3.11)
$(u^{k}_{t})\quad$ is bounded in $\quad L^{6}(Q).$ (3.12)
@@
$U^{k}\rightharpoonup U=(u,v)\quad\mbox{weakly* in}\quad L^{\infty}(0,T;\mathcal{H}).$ (3.13)
@@
From (3.11), Applying Lions’ compactness theorem [23, Ch. 1, Th. 5.1 ] with $H^{1}\hookrightarrow\hookrightarrow H^{0}\hookrightarrow H^{0}$, it turns out that there exists a subsequence of $(u^{k})$ still denoted by $(u^{k})$ such that
$u^{k}\rightarrow u\quad\mbox{strongly in}\quad H^{0}.$
Passing to a subsequence if necessary, we get
$u^{k}\rightarrow u\quad\mbox{a.e. in}\quad Q.$ (3.14)
Using the continuity of $f$, it follows from (3.14) that
$f(u^{k})\rightarrow f(u)\quad\mbox{a.e. in}\quad Q.$ (3.15)
@@
$f(u^{k})\quad\mbox{in bounded in}\quad L^{6/5}(Q).$ (3.16)
Then, using (3.15) and (3.16), it follows from Lions’ lemma [23, Ch. 1, Lemma 1.3] that
$f(u^{k})\rightarrow f(u)\quad\mbox{weakly in}\quad L^{6/5}(Q)=[L^{6}(Q)]^{\prime}.$ (3.17)
@@
$\lim\sup_{k,n\rightarrow+\infty}\int_{0}^{T}\int_{\Omega}\left[\,g(u_{t}^{k})-g(u_{t}^{n})\,\right]\left[\,u_{t}^{k}-u_{t}^{n}\,\right]dxdt\leq 0,$ (3.18)
@@
If $p<5$, then the last term in the above inequality is zero. However, for $p=5$ more delicate argument is needed. which depends on enhanced dissipation. To this end, we rewrite
$\begin{aligned} \int_{\Omega}\left[f(u^{k})-f(u^{n})\,\right]z^{k,n}_{t}dx=&\;\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta z^{k,n}z^{k,n}_{t}dx\\ =&\;\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta\frac{1}{2}\frac{d}{dt}|z^{k,n}|^{2}dx\\ =&\;\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\right]\\ &-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\,\right)d\theta|z^{k,n}|^{2}dx.\end{aligned}$ (3.21)
Then, substituting (3.21) in (3.19), we have
@@
$+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx=\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\,\right)d\theta|z^{k,n}|^{2}dx.$
@@
$f^{\prime}(s)\geq-\mu,\quad\mbox{for}\quad|s|>N.$ (3.23)
@@
$|f^{\prime}(s)|\leq C_{f}(1+|N|^{4}),\quad\mbox{for}\quad|s|\leq N.$ (3.24)
So, setting $\Omega=\Omega_{1}\cup\Omega_{2}$, where
$\Omega_{1}:=\{x\in\Omega:|\theta u^{k}+(1-\theta)u^{n}|>N\}\quad\mbox{and}\quad\Omega_{2}:=\{x\in\Omega:|\theta u^{k}+(1-\theta)u^{n}|\leq N\},$
it follows from (3.23) and (3.24) that
$\begin{aligned} &\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\\ &=\frac{1}{2}\left[\int_{\Omega_{1}}+\int_{\Omega_{2}}\right]\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\\ &\geq-\frac{\mu}{2}\int_{\Omega_{1}}|z^{k,n}|^{2}dx-\frac{C_{f}(1+|N|^{4})}{2}\int_{\Omega_{2}}|z^{k,n}|^{2}dx\geq-\frac{K_{f}}{2}\int_{\Omega}|z^{k,n}|^{2}dx,\end{aligned}$ (3.25)
where $K_{f}=\max\{\mu,C_{f}(1+|N|^{4})\}$.
On the other hand, from (2.4), Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and (3.11), we have
@@
$\begin{aligned} &\frac{d}{dt}E_{z^{k,n}}(T)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]|z^{k,n}|^{2}dx\\ &=\underbrace{K_{f}\int_{\Omega}z^{k,n}z^{k,n}_{t}dx}_{I_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\right)d\theta|z^{k,n}|^{2}dx}_{I_{2}}.\end{aligned}$ (3.28)
@@
$\left|I_{1}\right|\leq\frac{K_{f}}{\lambda_{1}^{1/2}}\|z^{k,n}_{t}(t)\|\|\nabla z^{k,n}(t)\|\leq\frac{K_{f}}{2\lambda_{1}^{1/2}}\left[\|z^{k,n}_{t}(t)\|^{2}+\|\nabla z^{k,n}(t)\|^{2}\right]\leq\frac{K_{f}}{\lambda_{1}^{1/2}}E_{z^{k,n}}(t).$ (3.29)
@@
$\begin{aligned} |I_{2}|\leq&\;C\left[\,1+\|u^{k}(t)\|_{6}^{3}+\|u^{m}(t)\|_{6}^{3}\,\right]\left[\,\|u^{k}_{t}(t)\|_{6}+\|u^{n}_{t}(t)\|_{6}\,\right]\|\|z^{n,k}(t)\|_{6}^{2}\\ \leq&\;C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t),\end{aligned}$ (3.30)
@@
$\frac{d}{dt}E_{z^{k,n}}(t)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx\leq C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t).$ (3.31)
@@
$\begin{aligned} \underbrace{E_{z^{k,n}}(t)}_{\geq 0}+\int_{0}^{t}\int_{\Omega}\underbrace{e^{C\int_{s}^{t}\psi(\tau,u^{k},u^{n})d\tau}}_{\geq 1}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dxds\leq e^{C\int_{0}^{t}\psi(s,u^{k},u^{n})ds}E_{z^{k,n}}(0),\end{aligned}$ (3.32)
@@
We note that the same conclusion as above holds under weaker assumptions imposed on the source $f$. It suffices that the constraints on the lower bound of $f;$ is imposed only for critical level of nonlinearity: $\lim_{|s|\rightarrow\infty}\frac{f^{\prime}(s)}{s^{4}}>-\nu$. The subcritical terms will converge to zero by the virtue of the compactness.
3.2 Proof of Theorem 2.2-(II): Energy Equality
@@
To proceed- let $v\in B(0,T;V)$ where $V$ is a Hilbert space and $B([0,T];V)$ is the space of $V$-value functions which are bounded on $[0,T]$, endowed with usual norm $\|y\|_{B([0,T];V)}=\sup_{t\in[0,T]}\|y(t)\|_{V}$. Let $\epsilon>0$ be a small parameter destined to go to zero and consider the extension $v(t)$ to all $t\in\mathbb{R}$ defined by
$v(t)=\left\{\begin{array}[]{l}v(T)\quad\mbox{if}\quad t\geq T,\\ v(0)\quad\mbox{if}\quad t\leq 0.\end{array}\right$.
With the above extensions we define
$v^{+}_{\epsilon}(t)$ $=$ $v(t+\epsilon)-v(t),$
$v^{-}_{\epsilon}(t)$ $=$ $v(t)-v(t-\epsilon),$
$D_{\epsilon}v(t)$ $=$ $\frac{1}{2\epsilon}\left[v^{+}_{\epsilon}(t)+v^{-}_{\epsilon}(t)\right].$
@@
    $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v(T)\|^{2}_{V}-\|v(0)\|^{2}_{V}\,\right].$ (3.35)
@@
    $\lim_{\epsilon\rightarrow 0}D_{\epsilon}v=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{+}_{\epsilon}=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{-}_{\epsilon}=v_{t}.$ (3.36)
@@
    $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v_{tt}(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v_{t}(T)\|^{2}_{V}-\|v_{t}(0)\|^{2}_{V}\,\right].$ (3.37)

### v4

@@ 2.2. Statements of the Main Results
In the case of subquintic sources, the analysis is simpler [S-S solutions are global], though still technical. Some of the recent relevant references include [27, 31] and references within, where nonlinear dissipation has been also treated. For the quintic waves with nonlinear dissipation some of the elements of the S-S method described above has been very recently applied to subquintic and quintic waves with nonlinear enhanced dissipation [33]. In particular, backward trajectories method patterned after [32] is the driving argument in establishing asymptotic compactness of the trajectories and their smoothness. This requires strong assumptions imposed on the damping [nondegenearcy of the slope at the origin] and also the source [coercivity of quintic estimate].
@@
One of the goals of this paper is to show that such restrictions, in general, are not needed. Moreover, in the nondegenerate case , quintic waves can be stabilised to finite dimensional sets. The model with enhanced dissipation takes its own life with no reference to either S-S solutions or backward trajectories. The main protagonist is the enhanced dissipation which could also be degenerate $g^{\prime}(0)=0$.] This alone provides asymptotic compactness effect for a large range of sources including the quintic-energy critical. No coercivity assumption is needed. The role of Strichartz estimates is taken, in some sense, by the enhanced dissipation generated by the nonlinearity of the damping. However, in the present case one is in a position to show that the enhanced dissipation is retained in the ultimate dissipativity estimate [unlike Strichartz estimate]. In view of this, the present work is a continuation of the model and methods considered [5] in the direction of dynamical systems and its long time behaviour. The main player in this game is energy equality established for weak solutions - an earlier version of the equality can be found in [5]. Adaptation of J.Ball’s method to nonlinear dissipation leads to the existence of global attractor. A general quasi-stability property for the dynamics is established under additional assumption of non-vanishing derivative of the damper. The latter implies finite-dimensionality, enhanced regularity of the attractor and also an existence of exponential attractor. In order to recover maximal regularity of the attractor, the splitting method used by Chueshov and Lasiecka in [14] in conjunction with gradient structure of the attractor and properties of the stationary set. In conclusion, the results provided eliminate critical restrictions in [33] and moreover lead to different methodology applicable to many other hyperbolic like dynamics with critical sources. The method is much simpler and it could also be applied to nonlocal damping under certain natural assumptions [33, 34].
@@
One should also mention works where strong damping is added to the model. But this is a form of regularization where originally hyperbolic dynamics becomes parabolic [linearisation generates an analytic semigroup]. The latter is a very different problem. References [11, 13, 19, 21].
Quintic waves with linear localised dissipation defined for stable dynamics has been considered in [12], where exponential decays of the energy are claimed.
Regular solutions for wave equation with super- critical [above quintic] exponents were considered in [8].
The remainder of this article is devoted to the proofs of the three theorems stated above. The definitions and results from the abstract theory of global attractors that are used herein are referred to the books [15, 16].
@@
3. Proof of Theorem 2.2: Weak Solutions and the Generation of the Dynamical System $(\mathcal{H},S_{t})$.
The well-posedness of the energy-critical wave equation with enhanced nonlinear dissipation has been investigated in [7]. Methods of monotone operator theory [3] were applied in order to construct semigroup solutions. However, several aspects of the theory presented in [7]-such as robustness and continuous dependence on the data, required additional hypotheses imposed on the data. These limitations are eliminated [necessary for construction of dynamical system] in the present approach which is based on construction of Galerkin solutions with the appropriate estimates exploring the enhanced dissipation. The estimates obtained are taken advantage of in the process of rigorous justification of several calculations throughout the paper..
@@
3.1. Proof of Theorem 2.2-(I): Construction of Weak Solutions
The proof of existence of weak solutions proceeds along general strategy used in [5], where the enhanced dissipation plays critical role in the arguments. However, the construction in [5] is based on suitable truncations of nonlinear terms and heavy use of monotone operator theory [3, 25]. In the present case, we resort to nonlinear Galerkin solutions, which also give a nice computational framework. The advantage of Galerkin solutions is that “formal” computations are well justified and lead to conserved quantities. However, critical limit passages on nonlinear terms rely on some arguments seen already in [5]. On the other hand, [5] deals also with the boundary sources of critical range.These will be more difficult to incorporate within the Galerkin scheme.
@@
Let $0<\lambda_{1}\leq\lambda_{2}\leq\cdots$ be the eigenvalues of the Laplacian operator $\Delta$ with Dirichlet boundary condition and $\omega_{1},\omega_{2},\cdots$ be the corresponding eigenfunctions such that they form an orthonormal basis in $H^{0}$. We assume the boundary $\Gamma=\partial\Omega$ regular enough such that $\omega_{j}\in H^{2}.$ Let $V_{k}=\mbox{Span}\{\omega_{1},\cdots,\omega_{k}\}$ the subspace of $H^{2}$ generated by the first $k$ elements of $\{\omega_{j}\}_{j\in\mathbb{N}}$. For $k\in\mathbb{N}$ we can construct a function $u^{k}$ given by $u^{k}(t)=\sum_{j=1}^{k}y_{jk}(t)\omega_{j}\in V_{k},$ $t\in[0,t_{k}),$ where $(y_{jk})$ is a local solution on $0,t_{k})\in[0,T)$ of the following system of ODEs:
(3.3) $\left\{\begin{array}[]{l}\displaystyle{(u^{k}_{tt}(t),\omega_{j})+(\nabla u^{k}(t),\nabla\omega_{j})+(g(u^{k}_{t}(t)),\omega_{j})+(f(u^{k}(t)),\omega_{j})=(h,\omega_{j})}\\ \displaystyle{U^{k}(0)=(u^{k}(0),u^{k}_{t}(0))=(u_{0k},u_{1k})=U_{0k},\quad j=1,\cdots,k.}\end{array}\right$.
@@
(3.4) $U^{k}(0)\longrightarrow U_{0}\quad\mbox{strongly in}\quad\mathcal{H}.$
Taking $\omega_{j}=u_{t}^{k}(t)$ in the approximate equation (3.3) and integrating over the interval $[0,t]$, we obtain that $u^{k}(t)$ satisfies the energy identity (2.8). Then, by the coercivity conditions given in (2.1) and (2.5), growth conditions (2.1) and (2.4), embeddings $H^{1}\hookrightarrow L^{6}(\Omega)\hookrightarrow H^{0}$, and convergence (3.4), we conclude that
@@
(3.6) $U^{k}=(u^{k},u^{k}_{t})\;\;\mbox{is bounded in}\;\;L^{\infty}(0,T;\mathcal{H});\;\;(u^{k}_{t})\;\;\mbox{is bounded in}\;\;L^{6}(Q).$
@@
(3.7) $U^{k}\rightharpoonup U=(u,v)\quad\mbox{weakly* in}\quad L^{\infty}(0,T;\mathcal{H}).$
@@
From (3.6), Applying Lions’ compactness theorem [26, Ch. 1, Th. 5.1 ] with $H^{1}\hookrightarrow\hookrightarrow H^{0}\hookrightarrow H^{0}$, it turns out that there exists a subsequence of $(u^{k})$ still denoted by $(u^{k})$ such that $u^{k}\rightarrow u$ strongly in $H^{0}.$ Passing to a subsequence if necessary, and by continuity of $f$ it follows:
(3.8) $u^{k}\rightarrow u\quad\mbox{a.e. in}\quad Q;~~f(u^{k})\rightarrow f(u)\quad\mbox{a.e. in}\quad Q.$
@@
(3.9) $f(u^{k})\quad\mbox{in bounded in}\quad L^{6/5}(Q).$
Then, using (3.8) and (3.9), it follows from Lions’ lemma [26, Ch. 1, Lemma 1.3] that
(3.10) $f(u^{k})\rightarrow f(u)\quad\mbox{weakly in}\quad L^{6/5}(Q)=[L^{6}(Q)]^{\prime}.$
@@
(3.11) $\lim\sup_{k,n\rightarrow+\infty}\int_{0}^{T}\int_{\Omega}\left[\,g(u_{t}^{k})-g(u_{t}^{n})\,\right]\left[\,u_{t}^{k}-u_{t}^{n}\,\right]dxdt\leq 0,$
@@
If $p<5$, then the last term in the above inequality is zero. However, for $p=5$ more delicate argument is needed which depends on enhanced dissipation. To this end, we rewrite
(3.14) $\begin{aligned} &\int_{\Omega}\left[f(u^{k})-f(u^{n})\,\right]z^{k,n}_{t}dx\\ &=\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\chi_{\theta})d\theta|z^{k,n}|^{2}dx\right]-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\chi_{\theta})(\chi_{\theta})_{t}d\theta|z^{k,n}|^{2}dx.\end{aligned}$
where $\chi_{\theta}:=\theta u^{k}+(1-\theta)u^{n}$. Then, substituting (3.14) in (3.12), we have
@@
$\quad+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx=\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\chi_{\theta})(\chi_{\theta})_{t}d\theta|z^{k,n}|^{2}dx.$
@@
(3.16) $f^{\prime}(s)\geq-\mu,\quad\mbox{for}\quad|s|>N.$
@@
(3.17) $|f^{\prime}(s)|\leq C_{f}(1+|N|^{4}),\quad\mbox{for}\quad|s|\leq N.$
So, setting $\Omega=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{1}:=\{x\in\Omega:|\chi_{\theta}|>N\}$ and $\Omega_{2}:=\{x\in\Omega:|\chi_{\theta}|\leq N\},$ it follows from (3.16) and (3.17) that
(3.18) $\begin{aligned} &\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime}(\chi_{\theta})d\theta|z^{k,n}|^{2}dx=\frac{1}{2}\left[\int_{\Omega_{1}}+\int_{\Omega_{2}}\right]\int_{0}^{1}f^{\prime}(\chi_{\theta})d\theta|z^{k,n}|^{2}dx\\ &\geq-\frac{\mu}{2}\int_{\Omega_{1}}|z^{k,n}|^{2}dx-\frac{C_{f}(1+|N|^{4})}{2}\int_{\Omega_{2}}|z^{k,n}|^{2}dx\geq-\frac{K_{f}}{2}\int_{\Omega}|z^{k,n}|^{2}dx,\end{aligned}$
where $K_{f}=\max\{\mu,C_{f}(1+|N|^{4})\}$. On the other hand, from (2.4), Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and (3.6), we have
@@
(3.21) $\begin{aligned} &\frac{d}{dt}E_{z^{k,n}}(T)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]|z^{k,n}|^{2}dx\\ &=\underbrace{K_{f}\int_{\Omega}z^{k,n}z^{k,n}_{t}dx}_{I_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\chi_{\theta})(\chi_{\theta})_{t}d\theta|z^{k,n}|^{2}dx}_{I_{2}}.\end{aligned}$
@@
(3.22) $\left|I_{1}\right|\leq\frac{K_{f}}{\lambda_{1}^{1/2}}\|z^{k,n}_{t}\|\|\nabla z^{k,n}\|\leq\frac{K_{f}}{2\lambda_{1}^{1/2}}\left[\|z^{k,n}_{t}\|^{2}+\|\nabla z^{k,n}\|^{2}\right]\leq\frac{K_{f}}{\lambda_{1}^{1/2}}E_{z^{k,n}}(t).$
@@
(3.23) $|I_{2}|\leq C\left[\,1+\|u^{k}\|_{6}^{3}+\|u^{m}\|_{6}^{3}\,\right]\left[\,\|u^{k}_{t}\|_{6}+\|u^{n}_{t}\|_{6}\,\right]\|\|z^{n,k}\|_{6}^{2}\leq C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t),$
@@
(3.24) $\frac{d}{dt}E_{z^{k,n}}(t)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx\leq C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t).$
@@
(3.25) $\underbrace{E_{z^{k,n}}(t)}_{\geq 0}+\int_{0}^{t}\int_{\Omega}\underbrace{e^{C\int_{s}^{t}\psi(\tau,u^{k},u^{n})d\tau}}_{\geq 1}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dxds\leq e^{C\int_{0}^{t}\psi(s,u^{k},u^{n})ds}E_{z^{k,n}}(0),$
@@
We note that the same conclusion as above holds under weaker assumptions imposed on the source $f$. It suffices that the constraints on the lower bound of $f$ is imposed only for critical level of nonlinearity: $\lim_{|s|\rightarrow\infty}\frac{f^{\prime}(s)}{s^{4}}>-\nu$. The subcritical terms will converge to zero by the virtue of the compactness.
3.2. Proof of Theorem 2.2-(II): Energy Equality
@@
To proceed- let $v\in B(0,T;V)$ where $V$ is a Hilbert space and $B([0,T];V)$ is the space of $V$-value functions which are bounded on $[0,T]$, endowed with usual norm $\|y\|_{B([0,T];V)}=\sup_{t\in[0,T]}\|y(t)\|_{V}$. Let $\epsilon>0$ be a small parameter destined to go to zero and consider the extension $v(t)$ to all $t\in\mathbb{R}$ defined by $v(t)=v(T)$ if $t\geq T$ and $v(t)=v(0)$ if $t\leq 0$. With these extensions, we define
$v^{+}_{\epsilon}(t)\equiv v(t+\epsilon)-v(t),~~v^{-}_{\epsilon}(t)\equiv v(t)-v(t-\epsilon),~~D_{\epsilon}v(t)\equiv\frac{1}{2\epsilon}\left[v^{+}_{\epsilon}(t)+v^{-}_{\epsilon}(t)\right].$
@@
    (3.27) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v(T)\|^{2}_{V}-\|v(0)\|^{2}_{V}\,\right].$
@@
    (3.28) $\lim_{\epsilon\rightarrow 0}D_{\epsilon}v=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{+}_{\epsilon}=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{-}_{\epsilon}=v_{t}.$
@@
    (3.29) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v_{tt}(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v_{t}(T)\|^{2}_{V}-\|v_{t}(0)\|^{2}_{V}\,\right].$